### v1

# PSDiffusion: Harmonized Multi-Layer Image Generation via Layout and Appearance Alignment

Abstract.
Diffusion models have made remarkable advancements in generating high-quality images from textual descriptions. Recent works like LayerDiffuse have extended the previous single-layer, unified image generation paradigm to transparent image layer generation. However, existing multi-layer generation methods fail to handle the interactions among multiple layers such as rational global layout, physics-plausible contact and visual effects like shadows and reflections while maintaining high alpha quality. To solve this problem, we propose PSDiffusion, a unified diffusion framework for simultaneous multi-layer text-to-image generation. Our model can automatically generate multi-layer images with one RGB background and multiple RGBA foregrounds through a single feed-forward process. Unlike existing methods that combine multiple tools for post-decomposition or generate layers sequentially and separately, our method introduces a global-layer interactive mechanism that generates layered-images concurrently and collaboratively, ensuring not only high quality and completeness for each layer, but also spatial and visual interactions among layers for global coherence.
Multi-Modal Generation, Diffusion Model, Layered Images
copyright: noneccs: Computing methodologies Image and video acquisition
Figure 1. Given a complex prompt, our model, PSDiffusion, produces plausible per-layer layouts for each instance and achieves harmonious inter-layer interactions. Compared to the state-of-the-art (Zhang and Agrawala, 2024), PSDiffusion generates more consistent appearances with realistic visual effects (e.g., shadows for the boat and swan), without manual refinement. [Best viewed with zoom-in]
1. Introduction
Diffusion-based generative models such as Stable Diffusion (Rombach et al., 2022) have revolutionized image synthesis by generating high-fidelity, diverse visuals from textual prompts, but they are limited to producing images as unified, single-layer outputs which are not easy to edit. In contrast, images with layered representations provide human designers with great advantages in image manipulations, such as precise editing through element isolation, compositional flexibility via layer recombination, and collaborative iteration through asset sharing.
The evolution of transparent image layer generation began with post-processing solutions applied to standard RGB outputs. Early works like PP-Matting (Chen et al., 2022) and Matting-Anything (Li et al., 2024) pioneered trimap-free alpha estimation for single transparent layer extraction. Subsequent pipelines such as MuLAn  (Tudosiu et al., 2024), PACO (Liu et al., 2024a) and LayeringDiff (Kang et al., 2025) established systematic frameworks for decomposing images into editable layers through detection, segmentation and completion. While operationally sound, these multi-stage methods tend to accumulate errors due to their sequential processing pipelines, primarily manifested in two aspects: challenges in generating high-quality alpha channels for foreground layers and inconsistencies arising from amodal completion operations.
This limitation catalyzed the emergence of direct RGBA synthesis paradigms. Text2Layer (Zhang et al., 2023b) marked a pivotal shift by jointly generating foreground/background layers through coordinated latent space optimization in Stable Diffusion. LayerDiffuse’s (Zhang and Agrawala, 2024) ”latent transparency” mechanism accelerated the field, which enabled pretrained LDMs to produce high-fidelity single transparent layers, and Alfie’s (Quattrini et al., [n. d.]) diffusion transformer architecture that dynamically derived alpha channels from attention maps. Building on transparent single-layer image generation, recent works have made strides toward multi-layer synthesis. LayerDiffuse (Zhang and Agrawala, 2024) itself extended its architecture for sequential background-to-foreground generation . Fontanella et al. (Fontanella et al., [n. d.]) explored independent multiple foreground layer generation and post-composition. LayerDiff (Huang et al., 2024) attempted to generate multi-layer images simultaneously, but it can only generate isolated, non-overlapping layers with binary masks. These paradigms mainly depend on straightforward layer stacking instead of considering multiple layers in a holistic manner. Therefore, they tend to neglect the interactions among layers, including comprehensive layout, physics-plausible contacts and visual effects like shadows and reflections, while these are essential to cohesive multi-layer generation and editing.
Another critical challenge underlying this mult-layer image generation stems from the scarcity of high-quality multi-layer RGBA image datasets. While text-image pairs abound in datasets like LAION (Schuhmann et al., 2021), commercially restricted access to premium transparent assets severely limits usable multi-layer RGBA data. Existing open-source multi-layer datasets exhibit complementary weaknesses. MAGICK (Burgert et al., 2024) and Multi-layer Dataset (Zhang and Agrawala, 2024) provide visually coherent layers but are constrained to 1-2 layers per composition, whereas MuLAn (Tudosiu et al., 2024) supports 2-6 layers per image but suffers from poor alpha-matte quality because it relies on automated post-processing pipelines.
To overcome these dataset limitations, we propose the Inter-Layer Dataset, which consists of 30K meticulously curated multi-layer images, each containing 3-6 layers with artist-grade alpha mattes and rich layer interactions. Recognizing that automated synthesis and post-editing inherently fail to guarantee precise alpha channels and contextually coherent inpainting, we engaged professional designers to construct this dataset through a human-in-the-loop workflow. Designers curated layers from diverse sources including web repositories, BG20K (Li et al., 2022), MAGICK (Burgert et al., 2024), and LayerDiffuse-generated transparent assets (Zhang and Agrawala, 2024), and then applied meticulous edits to ensure natural layer interactions. Each composition integrates spatial layout optimization, physics-aware interaction and global stylistic harmony.
Although existing pretrained text-to-image generation models can only generate unified single layer images, given a global text prompt describing the background, multiple foregrounds and their interactions, they did demonstrate prominent capability to spontaneously arrange the layout of multiple elements, ensure natural interactions between constituent subjects, and maintain global harmony and consistency in generated images. We believe that these pretrained text-to-image generation can provide interaction priors for multi-layer generation. Building on this fact, we propose PSDiffusion, a unified diffusion framework that utilizes the interaction priors from pretrained diffusion models, to simultaneously generate multi-layer images with coherent global structure and realistic layer interactions. We design an attention-based global-layer interactive mechanism to model the layer interactions paradigm. For one thing, to achieve a harmonious global layout for foreground layers, we develop a layer cross-attention reweighting module, where we extract the cross-attention map from global layer to guide the position of foreground layers. For another, to facilitate content coherence and interaction among different layers, and avoid layer entanglement, we develop a partial joint self-attention module, where we encourage content sharing from global layer across the foreground and background layers.
Given a global prompt and various layer-specific prompts as input, our PSDiffusion model enables simultaneous multi-layer RGBA image generation with enhanced interactions among all layers. The layer-specific prompts can either be automatically decomposed by LLMs or manually defined by users for precise layer control. Furthermore, with the layered image representations, our model supports various layer-wise editing for users to implement precise layer control. In summary, our contributions are as follows:
  • We propose Inter-Layer, a high quality multi-layer RGBA image dataset consisting of 30K samples, each containing 3-6 layers with artist-grade alpha mattes and rich, harmonious layer interactions.
  • We present PSDiffusion, an end-to-end diffusion framework for simultaneous multi-layer image generation. Leveraging a global-layer interactive mechanism, our model generates layered images concurrently and collaboratively, ensuring not only high quality and completeness for each layer, but also spatial and visual interactions among layers for global coherence.
2. Related Works
2.1. Text-to-Image Generation
Diffusion models have made significant advancements in image generation, particularly in text-guided image synthesis, leading to remarkable improvements in both the quality of generated images and the ability to follow complex prompts. The development of large-scale diffusion models trained on extensive datasets of paired text and images has established new benchmarks in this domain, with notable examples including Stable Diffusion (Rombach et al., 2022), SDXL (Podell et al., 2024), Imagen (Saharia et al., 2022), and DALL-E 3 (Betker et al., 2023). Furthermore, recent advancements in network architecture, especially with Diffusion Transformer networks (DiT) like Stable Diffusion 3 (Esser et al., 2024) and Flux (Labs, 2024) , have enhanced image fidelity and diversity. Moreover, other research efforts focus on integrating various control conditions into pre-trained image foundation models to enhance their controllability. For instance, Gligen (Li et al., 2023), ControlNet (Zhang et al., 2023a) and IP-Adapter (Ye et al., 2023) offer greater flexibility in controllable image generation, allowing users to precisely manipulate aspects such as layout, depth, human pose, and style. However, these models can only generate a single RGB image during a single forward pass, rather than producing layered images, which are widely used formats that enable users to edit specific parts of an image without affecting other areas.
2.2. Layered Image Generation
Image Layer Extraction with Post-RGB Process. Multi-layer image generation can be achieved primarily through two different paths. The first path, termed Post-RGB Process, operates through a sequential pipeline: (1) initial RGB image synthesis via diffusion models (Rombach et al., 2022); (2) foreground extraction using open-vocabulary object detection and segmentation tools (e.g. GroundingDINO (Liu et al., 2024b), SAM (Kirillov et al., 2023), Matting-Anything (Li et al., 2024)); (3) background inpainting and foreground completion for occluded areas to derive multiple RGBA layers (Rombach et al., 2022; Podell et al., 2024; Yu et al., 2023; Ao et al., 2024). Along this way, Tudosiuet et al. (Tudosiu et al., 2024) developed a training-free pipeline, which decomposes a monocular RGB image into a stack of RGBA layers comprising background and isolated instances, and released the MuLAn dataset. LayeringDiff  (Kang et al., 2025) begins by generating using an off-the-shelf image generative model, followed by disassembling the image into its constituent foreground and background layers. Yang et al. proposed LayerDecomp (Yang et al., 2024), a generative training framework that leverages both simulated and real-world data to enable robust layer-wise decomposition with accurate visual effect representation through a consistency loss, facilitating high-quality, training-free downstream editing. Generative Omnimatte (Lee et al., 2024) extends this idea to decompose video into layers and demonstrate high-quality decompositions and editing results. Despite the operational validity of these methods, the multi-stage post-processing approach imposes a nontrivial computational redundancy and risks compounding errors through iterative operations—a limitation that frequently manifests itself as incoherent visual semantics or fragmented stylistic attributes.
Direct Transparent Image Layer Generation. RGBA direct generation focuses on synthesizing image layers directly, bypassing the traditional pipeline of generating a full image first and then extracting transparent layers. The earliest work, Text2Layer (Zhang et al., 2023b), extends the Stable Diffusion architecture into a dual-layer generation framework. It jointly synthesizes a background scene and semantically aligned foreground elements through coordinated latent space optimization. Subsequently, LayerDiffuse (Zhang and Agrawala, 2024) introduces ”latent transparency”, enabling large-scale pretrained latent diffusion models to generate either single transparent images. This approach has gained a reputation for its high-quality single-layer transparency generation. In parallel, Alfie (Quattrini et al., [n. d.]) and Diffumatting (Hu et al., 2024a) also develop different pipelines to generate single-layer RGBA images.
Building on single-layer generation capabilities, multi-layer synthesis is achieved through either sequential or separate generation paradigms. Specifically, LayerDiffuse (Zhang and Agrawala, 2024) adopts a sequential approach, proposing a background-conditioned transparent layer generation that iteratively synthesizes layers from background to foreground. Fontanella et al. (Fontanella et al., [n. d.]) and LayerFusion (Dalva et al., 2024) employ a separate paradigm, where individual foreground layers are generated independently and then composited.
While methods of multi-layer generation based on single-layer generation have advanced RGBA synthesis, their extension suffers from insufficient layer interaction, leading to critical practical limitations. Common inter-layer contamination issues include rational global layout, physics-plausible contacts and visual effects like shadows and reflections. These problems are not easily resolvable through additional post-editing, particularly for the latter two issues.
Recent unified text-guided multi-layer frameworks have emerged and shown promising results. Specifically, LayerDiff (Huang et al., 2024) proposes a simultaneous multi-layer synthesis framework using a unified diffusion model. However, it produces low-quality outputs limited to binary masks instead of full alpha channels, failing to handle overlapping foregrounds and visual effects which are essential for layer editing. ART (Pu et al., 2025) introduces an Anonymous Region Layout Planner, aligning visual tokens with text tokens to directly synthesize spatially coherent multi-layer images. While innovative, it relies on precise bounding boxes provided by users or generated by LLMs. Additionally, its anonymous layout mechanism results in layer entanglement issue, where semantically distinct elements (e.g., foreground characters and background props) are incorrectly merged into one layer, adding unnecessary complexity.
In contrast, our proposed PSDiffusion is a unified diffusion framework for simultaneous text to multi-layer image generation, ensuring not only high quality and completeness for each layer, but also spatial and semantic interaction among layers for global coherence. We adopt global prompt and layer-specific prompts as input to facilitate customized element separation and avoid layer entanglement.
Figure 2. Architecture of PSDiffusion. Our framework processes multi-layer compositions through three key components: (1) A transparent VAE encoder preserving alpha channels; (2) Layer cross-attention reweighting for layout plausibility; (3) Partial joint self-attention for inter-layer context modeling.
3. Method
3.1. Problem Formulation
A multi-layer composable image is defined as a structure containing a background layer $B=I_{c}^{B}\in\mathbb{R}^{H\times W\times 3}$ (RGB format) and $K$ foreground layers $\{F_{k}\}_{k=1}^{K}$ in RGBA format, where each $F_{k}=\left(I_{c}^{F_{k}},I_{\alpha}^{F_{k}}\right)$ consists of color channels $I_{c}^{F_{k}}\in\mathbb{R}^{H\times W\times 3}$ and an alpha channel $I_{\alpha}^{F_{k}}\in\mathbb{R}^{H\times W\times 1}$. Following the premultiplied alpha convention, each foreground is preprocessed as $F^{\prime}_{k}=I_{c}^{F_{k}}\ast I_{\alpha}^{F_{k}}$ through element-wise multiplication.
The composition pipeline initializes with the background $\smash{I_{0}=I_{c}^{B}}$, then recursively blends layers via alpha operations. The $k$-th composite image is given by
(1) $I_{k}=\left(1-I_{\alpha}^{F_{k}}\right)\cdot I_{k-1}+I_{\alpha}^{F_{k}}\ast I_{c}^{F_{k}},\ k\in[K]$
where $F_{k}$ denotes the $k$-th foreground. After $K$ iterations, the final image $I_{K}$ composes into
(2) $I_{K}=\underbrace{\left(\prod_{k=1}^{K}\left(1-I_{\alpha}^{F_{k}}\right)\right)\ast I_{c}^{B}}_{\text{\scriptsize Background visibility attenuated}}+\underbrace{\sum_{k=1}^{K}\left(\prod_{j=k+1}^{K}\left(1-I_{\alpha}^{F_{j}}\right)\ast I_{\alpha}^{F_{k}}\ast I_{c}^{F_{k}}\right)}_{\text{\scriptsize Foreground contributions with occlusion}}$
This mathematical formulation explicitly captures two fundamental mechanisms: (1) The multiplicative attenuation of background visibility through successive transparency operations; (2) The nested modulation of foreground layer contributions based on depth-ordered occlusions.
3.2. PSDiffusion
Existing multi-layer image generation models contend with inherent challenges in synthesizing coherent compositions and fostering harmonious interactions between distinct high-quality transparent foreground layers and the background image (Tudosiu et al., 2024; Zhang and Agrawala, 2024). However, it has been observed that pretrained image foundation models demonstrate proficiency in generating intricate and visually coherent compositions, effectively integrating various objects and backgrounds within a single RGB image, owing to their training on extensive datasets derived from real-world imagery. Our primary insight is to exploit the layout and interaction priors from the denoising process of the global RGB image, to inform the generation of transparent foreground layers and background layer.
To fully take advantage of the layout arrangement and appearance harmonization of the global RGB images, we propose a novel framework, denoted as PSDiffusion, which concurrently generates multiple layers in a single feedforward pass, ensuring both reasonable layouts and harmonized interactions among the layers. As delineated in Figure 2, PSDiffusion comprises three principal components. First, PSDiffusion finetunes a pretrained RGB image Variational Autoencoder (VAE) into an RGBA VAE, thereby facilitating the generation of the alpha channel requisite for RGBA images, as articulated by LayerDiffuse (Zhang and Agrawala, 2024). Second, we introduce a layer cross-attention reweighting mechanism that extracts spatial layout information from the global image containing comprehensive content, generated via a global prompting strategy. This design paradigm ensures a coherent arrangement of disparate foreground objects. Finally, to promote harmonious interactions among the various foreground layers, we implement a partial joint self-attention module for inter-layer context modeling. Detailed descriptions of these two modules are provided as follows.
Figure 3. The overview of layer cross-attention reweighting module and partial joint self-attention mechanism.
Layer Cross-Attention Reweighting. To ensure a reasonable layout arrangement of foreground layers, we propose a novel approach that diverges from recursive layer generation. Our key insight is to refer to the layout arrangement from the denoising process of a complete RGB image generated using the global prompt. As illustrated in Figure 3, we utilize the cross-attention block of pre-trained image foundation model to extract the layout priors. Leveraging the emergent correspondence between prompts and features—a phenomenon widely demonstrated and utilized in existing literature (Hertz et al., [n. d.]; Tang et al., 2023; Cao et al., 2023; Epstein et al., 2023; Hu et al., 2024b)—we can precisely derive and control the location of foreground instance by calculating and manipulating the cross-attention map associated with the foreground instance token. As is highlighted in the prompts in Figure 3, we select the token of the noun representing the foreground subject because its attention map most prominently displays the location of the foreground subject. Specifically, at timestep $t$ in the $n$-th cross-attention block, the global text embedding $\psi(p)$ is projected to a key matrix $\mathbf{K}^{n}=\ell_{\mathbf{K}}^{n}(\psi(p))$ and a value matrix $\mathbf{V}^{n}=\ell_{\mathbf{V}}^{n}(\psi(p))$ , and the deep spatial features of the noisy image $\phi(z_{t})$ is projected to a query matrix $\mathbf{Q}^{n}=\ell_{\mathbf{Q}}^{n}(\phi(z^{t}))$, via learned linear projections $\ell_{\mathbf{K}}^{n}$, $\ell_{\mathbf{V}}^{n}$, and $\ell_{\mathbf{Q}}^{n}$, respectively. The global attention map $\mathbf{M}^{\mathrm{g}}\in\mathbb{R}^{hw\times s}$can be calculated by
(3) $\mathbf{M}^{\mathrm{g}}=\mathrm{Softmax}\left(\dfrac{\mathbf{Q}^{n}\left(\mathbf{K}^{n}\right)^{\mathrm{T}}}{\sqrt{d}}\right).$
We extend our method to obtain the layer attention map $M^{l}\in\mathbb{R}^{hw\times s}$ in the latent feature of the transparent image layer using an analogous calculation approach, where $hw$ denotes the size of the latent feature and $s$ denotes the length of the tokens. We extract the specific attention-map of the token representing the foreground subject from the global map and layer map, denoted as $\mathbf{M}_{i}^{\mathrm{g}}\in\mathbb{R}^{hw}$ and $\mathbf{M}_{i}^{\mathrm{l}}\in\mathbb{R}^{hw}$. Then implement the layer cross attention map reweighting by
(4) ${\tilde{\mathbf{M}}_{i}^{l}}$ $=\mathrm{Norm}\left(\mathbf{M}_{i}^{\mathrm{g}}\right)*\mathbf{M}_{i}^{l},$
(5) ${\hat{\mathbf{M}}_{i}^{l}}$ $={\tilde{\mathbf{M}}_{i}^{l}}/\mathrm{max}\left(\tilde{\mathbf{M}}_{i}^{l}\right)*\mathrm{max}\left(\mathbf{M}_{i}^{l}\right),$
where $\mathrm{Norm}(\cdot)$ denotes Min-Max normalization on the whole feature dimension and $\max(\cdot)$ is applied on the feature dimension as well. This process provides a more precise attention map that clearly delineates the position and shape of areas associated with the foreground layer prompt. The refined attention mask is then strategically integrated into the self-attention blocks of the latent diffusion model. Through this approach, we achieve a more nuanced and controlled layout arrangement for multi-layer image generation.
Partial Joint Self-attention. While our proposed layer cross-attention reweighting mechanism ensures a reasonable layout arrangement for each layer, challenges remain in harmonizing the color, style and interactions across different layers. To address the appearance inconsistency among layers, we introduce a novel partial joint self-attention mechanism for cross-layer feature context modeling. This approach facilitates more nuanced interactions among layers, ultimately achieving consistent color and lighting conditions within the composite image. As illustrated in Figure 3, we concatenate the features from global branch into the $n$-th self-attention block of the layer branch,
(6) $\mathbf{Q}=\mathbf{Q}_{L}=\mathbf{W}^{n}_{\mathbf{Q}}z_{i},\ \mathbf{K}=\mathbf{W}^{n}_{\mathbf{K}}\left[\mathbf{K}_{GL},\mathbf{K}_{L}\right],\ \mathbf{V}=\mathbf{W}^{n}_{\mathbf{V}}\left[\mathbf{V}_{GL},\mathbf{V}_{L}\right],$
where $z_{i}$ denotes the latent feature for the $i$-th transparent layer, [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes the concatenation operation. $\mathbf{K}_{GL}$ and $\mathbf{V}_{GL}$ is calculated by
(7) $\mathbf{K}_{GL}=\{\mathbf{K}_{G}\cdot\mathrm{Mask}_{i}^{\mathrm{g}}\},\ \mathbf{V}_{GL}=\{\mathbf{V}_{G}\cdot\mathrm{Mask}_{i}^{\mathrm{g}}\}.$
[…]
Overall, such a design enables the model to refer the appearance of the global layer, promoting a more cohesive and harmonized image generation across different foreground transparent layers and background layer.
Figure 4. Examples of our proposed inter-layer dataset.
3.3. Inter-Layer Dataset
To address the limitations of existing multi-layer datasets, we developed the dataset Inter-Layer through a human-centric curation pipeline. This dataset comprises 30,000 high-fidelity multi-layer compositions, each containing 3 - 6 layers with professional-grade alpha mattes and sophisticated inter-layer relationships. Recognizing the inadequacy of automated synthesis for achieving precise transparency channels and contextual coherence, we engaged professional designers in a multi-stage workflow. Curators systematically aggregated assets from diverse sources including web repositories, BG20K (Li et al., 2022), MAGICK (Burgert et al., 2024), and LayerDiffuse-generated transparent assets (Zhang and Agrawala, 2024), followed by rigorous manual refinement. Designers optimized each composition through: (1) Spatial layout organization with an aesthetic consideration; (2) Physics-plausible interactions among all layers; (3) Global visual harmonization across all elements. This human-in-the-loop approach ensures both technical precision in alpha channel quality and visual consistency in multi-layer storytelling, significantly advancing beyond existing datasets’ capabilities in layer complexity and visual authenticity. Some figures in our dataset are shown in Figure 4.
Figure 5. Qualitative comparison with state-of-the-art methods. Compared to existing methods, our model synthesizes multi-layer images with reasonable layout arrangement, harmonious appearance with visual effects and high alpha quality. Zoom in to check the details of foreground and composite images like shadows and edges. As exemplified by the multi-layer ensemble in the upper-right quadrant, the shadows of the elephant, man and woman are consistent with the background in their respective positions.
4. Experiments
4.1. Implementation Details
To better preserve the interaction priors of existing models, PSDiffusion is fine-tuned from Stable Diffusion XL (Podell et al., 2024) using LoRA (Hu et al., 2022). For global layer, we trained the additional $\mathbf{K}_{GL}$,$\mathbf{V}_{GL}$ linear layers in all self-attention blocks. For foreground layers and background layers, we trained $\mathbf{Q}$,$\mathbf{K}$,$\mathbf{V}$,$\mathbf{Out}$ linear layers in all attention blocks. Layer Cross-Attention Reweighting is implemented in layers at the resolution of 16, and the Partial Joint Self-attention is applied to all self-attention layers. We use the pre-trained VAE from Stable Diffusion XL for the global and background branch, and fine-tune it into an RGBA VAE to support the alpha channel of RGBA foreground image following LayerDiffuse (Zhang and Agrawala, 2024). During training, all layers of a multi-layer sample are encoded into latent space and added with the same noise at the same timestep. We use the AdamW optimizer and set the learning rate at $-41e\text{-}$ 4. Training is conducted with a total batch size of 8 on 20 NVIDIA A800 GPU hours for 10,000 iterations.
4.2. Compared with State-of-the-Art
Composite Image Layer Image
[…]
LayerDiffuse  (Zhang and Agrawala, 2024) 29.77 89.12 0.265 0.214 31.25 85.85
ART  (Pu et al., 2025) 29.93 96.54 0.743 0.702 128.1
PSDiffusion(ours) 31.89 87.32 0.766 0.751 31.76 83.71
Table 1. Quantitative comparison with state-of-the-art methods. We report performance on both composite images and layer images in terms of visual quality, text alignment and composite layout harmony and interaction plausibility.
Figure 6. More visual results of PSDiffusion. Through the design of Layer Cross-Attention Reweighting Module and Partial Joint Self-Attention Mechanism, PSDiffusion synthesize multi-layer transparent images with plausible layout arrangements and harmonious appearance.
We compare the quality of multi-layer images generated by our model with LayerDiffuse (Zhang and Agrawala, 2024) and the latest work ART (Pu et al., 2025), as they are the state-of-the-art layered image synthesis methods, and the only methods that release the source code. For evaluation, we construct 5,000 sets of multi-layer prompts, each composed of three foreground prompts, one background prompt and a global prompt describing their spatial and semantic interactions. We use vision-language model (Bai et al., 2025) to generate the prompts to establish diverse evaluation scenarios.
We conduct quantitative evaluation using multiple metrics. For image quality assessment, we measure feature distribution discrepancies between synthesized layers and reference datasets via FID (Heusel et al., 2017) (Fréchet Inception Distance). Specifically, the RGBA layer is evaluated against the MAGICK dataset (Burgert et al., 2024), while the composite images are compared with the COCO dataset (Lin et al., 2014). To assess text-image alignment for each layer, we employ the CLIP Score metric (Hessel et al., 2021), which quantifies the cosine similarity between image features and text prompts. To evaluate the harmony of the global layout and the plausibility of the layer interactions, we utilize vision-language model (Bai et al., 2025) to conduct scoring.
Qualitative Results. In Figure  5, we present the qualitative results on text prompt multi-layer image generation. LayerDiffuse (Zhang and Agrawala, 2024) tends to stacks all layers at the center of the image, neglecting the spatial interactions among all layers. The latest work ART showcases reasonable layout but relies on the precise bounding boxes given by users. It also exihibits the problem of imperfect alpha quality and layer entanglement. In contrast, our model displaces ont only high quality RGBA layers, but also the capability to arrange the layers to reasonable scale and postion, and mainting the visual interactions among the layers like the semi-transparent shadows. Zoom in to see the details of foreground and composite images like edges and shadows.
Quantitative Results. We report quantitative results in Table  1. We compare our PSDiffusionwith the state-of-the-art multi-layer image generation models LayerDiffuse (Zhang and Agrawala, 2024) and ART (Pu et al., 2025) on composite images. For LayerDiffuse (Zhang and Agrawala, 2024), we implement the proposed sequential background-to-foreground geneartion to gain the multi-layer images. For ART (Pu et al., 2025), since it relys on the bounding-boxes as input, we first apply the LLM layout generation proposed by the authors given our global prompt, and then use the generated bounding boxes and global prompts to generate the multi-layer images. As shown in the table, our PSDiffusion outperforms LayerDiffuse and ART across four metrics. The results on layout harmony and interaction plausibility directly showcase our strengths on generating multi-layer image with enhanced layer interaction and global rationality. Furthermore, for the alpha image layer, the results in Table  1 shows that our PSDiffusion also achieves excellent alpha image quality and better text alignment.
User Study. A comprehensive perceptual evaluation was conducted with 30 participants recruited from academic institutions to holistically evaluate the visual synthesis performance of multi-layer image generation models. The study encompassed totally 171 visual samples derived from 15 distinct multi-layer text prompts, with each case containing three components: foreground layers, backgrounds layers and composite outputs generated by two baseline methods alongside our proposed PSDiffusion. Each sample underwent multi-criteria evaluation using 5-point Likert scales (1: deficient, 5: exceptional) focusing on three critical dimensions: (1) semantic-textual congruence measuring prompt relevance and contextual appropriateness, (2) visual-perceptual fidelity evaluating aesthetic coherence and artifact suppression, and (3) structural clarity assessing image definition and detail preservation. The comparative evaluation results are summarized in Table  2, demonstrating that our method consistently achieves higher user ratings across than other methods. Complete experimental details regarding the user study design and implementation are available in the supplementary materials.
[…]
LayerDiffuse (Zhang and Agrawala, 2024) 1.64 3.63 1.53
ART (Pu et al., 2025) 1.30 2.53 3.14
PSDiffusion(ours) 2.06 3.83 3.35
Table 2. User study on multi-layer image generation models. Results generated by our PSDiffusionare the most popular with users.
4.3. Ablation studies
[…]
w/o Reweighting 31.64 89.18 0.565 0.418
w/o Joint Attention 30.71 89.54 0.488 0.350
Full model 31.89 87.32 0.766 0.751
Table 3. Ablation study on Layer Cross-Attention Reweighting and Partial Joint Self-Attention on composite images
Layer Cross-Attention Reweighting. In our layer cross-attention reweighting, we utilize the cross-attention map of the global layer, and implement the reweighting operation on the cross attention map of the foreground layers to control the global layout. As shown in Figure 7, without the layer attention-map reweighting, the foreground object tends to occupy the most of the image like LayerDiffuse and even shows poor quality on the alpha channel, especially for the foreground that should have been small in the composite image.
Figure 7. Visual results of ablation study for layer cross-attention reweighting. Full PSDiffusion exihibits more plausible spatial layout including postion and scale.
Partial Joint Self-Attention. In our partial joint self-attention blocks, we implement the content sharing from global layer across foreground and background layers to encourage layer interaction plausibility like the shadows. Figure  8 displays that removing this block leads to visual effects like rigid stacks of separated foreground layers hence global disharmony. As shown in Table  3, removing either Partial joint self-attention block or Layer cross-attention reweighting block results in significant decline in both global harmony and layer interactions. This showcases their entangled importance to the coherence and interaction plausibility of the composite image.
Figure 8. Visual results of ablation study for partial joint self-attention. Full PSDiffusion demonstrates more coherent appearances and visual effects (e.g., shadows).
5. Conclusion
In this paper, we propose the Inter-Layer Dataset, a high-quality multi-layer RGBA image dataset with artist-grade alpha mattes and rich layer interactions. Building on this, we introduce PSDiffusion, an end-to-end diffusion framework for simultaneous multi-layer image generation. We design a global-layer interactive mechanism to extract interaction priors from pre-trained diffusion models. Our model generates layered images concurrently and collaboratively, ensuring not only high quality and completeness for each layer, but also spatial and visual interactions among layers for global coherence.
For one thing, to achieve a harmonious global layout for foreground layers, we develop a layer cross-attention reweighting module, which extracts the cross-attention maps from the global branch to guide the positioning of foreground layers. For another, to enable content coherence and interaction among different layers, and avoid layer entanglement, we develop a partial joint self-attention module, where we facilitate content sharing from global layer across the foreground and background layers. With these layout and appearance alignment, we empower diffusion models to generate multi-layer images with enhanced layer interactions and composite harmony.
Supplementary Materials
Appendix A Limitations
A.1. Planar composition constraints in volumetric interaction modeling
While our model achieves sophisticated occlusion handling through discrete 2D layer composition, its planar representation inherently limits the modeling of volumetric spatial interactions. The strict layer-wise depth prioritization (e.g., globally enforcing Layer A to occlude Layer B) cannot capture localized mutual occlusion in 3D space, such as intertwined limbs during human embraces. Static layer ordering further fails to adapt to perspective-dependent depth transitions under viewpoint changes, while non-planar intersections with depth-dependent partial visibility remain unresolved due to reliance on discrete alpha compositing. These constraints stem from our use of discretized depth ordering rather than continuous 3D occupancy field learning.
A.2. Uniform sharpness artifacts from 2D-flattened rendering
Despite achieving exceptional per-layer clarity and lighting consistency, the uniform maximum sharpness across all layers prevents emulation of natural depth-of-field effects (e.g., foreground-background focus gradients). This limitation arises from the absence of depth-aware dynamic blurring, compromising visual realism in scenes requiring strong 3D spatial perception. The persistent all-in-focus appearance introduces planar artifacts, particularly noticeable in multi-layer compositions with complex depth relationships, as the discrete compositing mechanism cannot simulate depth-dependent optical blur.
Appendix B Dataset Visualization
Benefiting from our human-centric curation pipeline and professional-grade multi-stage refinement, our dataset achieves three key superior qualities: high-fidelity alpha mattes, physics-aware inter-layer interactions, and globally harmonized visual aesthetics. Fig. 910 are visualizations of our dataset compositions.
Figure 9. Visualization of our multi-layer dataset
Figure 10. More visulization of our multi-layer dataset
Appendix C User study
The following Table 4 presents the questionnaire form we made for the investigation in the userstudy part.
Table 4. User Study of Multi-Layer Generation Models
Current AI image generators create flat images, unlike professional design workflows using layered files (e.g., Photoshop .psd). Our new system automatically generates:
multi-layer images (1 background + multiple RGBA layers) Context-aware layer coordination Production-ready layered files
Evaluation Criteria
1. Semantic-textual congruence
Relevance to prompts & contextual logic Visual-perceptual fidelity 2. Aesthetic unity & artifact-free rendering
3. Structural clarity
Detail preservation & layer definition
Your Task
For each text-generated image:
Assign a 1-5 overall score (1=worst, 5=best) Base your rating on the three criteria above
Example Case
”whole text”: A white cat sitting and looking at the camera, positioned in front of a wooden floor Foreground:A white cat [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Score Background:wooden floor [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Score Merged image [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Score

### v2

@@
 PSDiffusion: Harmonized Multi-Layer Image Generation via Layout and Appearance Alignment
-Abstract.
-Diffusion models have made remarkable advancements in generating high-quality images from textual descriptions.
-Recent works like LayerDiffuse have extended the previous single-layer, unified image generation paradigm to transparent image layer generation.
-However, existing multi-layer generation methods fail to handle the interactions among multiple layers such as rational global layout, physics-plausible contact and visual effects like shadows and reflections while maintaining high alpha quality.
-To solve this problem, we propose PSDiffusion, a unified diffusion framework for simultaneous multi-layer text-to-image generation.
-Our model can automatically generate multi-layer images with one RGB background and multiple RGBA foregrounds through a single feed-forward process.
-Unlike existing methods that combine multiple tools for post-decomposition or generate layers sequentially and separately, our method introduces a global-layer interactive mechanism that generates layered-images concurrently and collaboratively, ensuring not only high quality and completeness for each layer, but also spatial and visual interactions among layers for global coherence.
-Multi-Modal Generation, Diffusion Model, Layered Images
-copyright: noneccs: Computing methodologies Image and video acquisition
-Figure 1.
-Given a complex prompt, our model, PSDiffusion, produces plausible per-layer layouts for each instance and achieves harmonious inter-layer interactions.
-Compared to the state-of-the-art (Zhang and Agrawala, 2024), PSDiffusion generates more consistent appearances with realistic visual effects (e.g., shadows for the boat and swan), without manual refinement.
+Abstract
+Transparent image layer generation plays a significant role in digital art and design workflows.
+Existing methods typically decompose transparent layers from a single RGB image using a set of tools or generate multiple transparent layers sequentially.
+Despite some promising results, these methods often limit their ability to model global layout, physically plausible interactions, and visual effects such as shadows and reflections with high alpha quality due to limited shared global context among layers.
+To address this issue, we propose PSDiffusion, a unified diffusion framework that leverages image composition priors from pre-trained image diffusion model for simultaneous multi-layer text-to-image generation.
+Specifically, our method introduces a global layer interaction mechanism to generate layered images collaboratively, ensuring both individual layer quality and coherent spatial and visual relationships across layers.
+We include extensive experiments on benchmark datasets to demonstrate that PSDiffusion is able to outperform existing methods in generating multi-layer images with plausible structure and enhanced visual fidelity.
+Figure 1: Given a complex prompt, our model, PSDiffusion, produces plausible per-layer layouts for each instance and achieves harmonious inter-layer interactions.
+Compared to the state-of-the-art [zhang2024transparent], PSDiffusion generates layered images with more plausible spatial arrangement and more consistent appearances with realistic visual effects (e.g., reflections for the boat and swan), without manual spatial refinement.
 [Best viewed with zoom-in]
-1. Introduction
-Diffusion-based generative models such as Stable Diffusion (Rombach et al., 2022) have revolutionized image synthesis by generating high-fidelity, diverse visuals from textual prompts, but they are limited to producing images as unified, single-layer outputs which are not easy to edit.
+1 Introduction
+Diffusion models [rombach2022high, dalle3, imagen] have revolutionized image synthesis by generating high-fidelity, diverse visuals from textual prompts, but they are limited to producing images as unified, single-layer outputs which are not easy to edit.
 In contrast, images with layered representations provide human designers with great advantages in image manipulations, such as precise editing through element isolation, compositional flexibility via layer recombination, and collaborative iteration through asset sharing.
-The evolution of transparent image layer generation began with post-processing solutions applied to standard RGB outputs.
-Early works like PP-Matting (Chen et al., 2022) and Matting-Anything (Li et al., 2024) pioneered trimap-free alpha estimation for single transparent layer extraction.
-Subsequent pipelines such as MuLAn  (Tudosiu et al., 2024), PACO (Liu et al., 2024a) and LayeringDiff (Kang et al., 2025) established systematic frameworks for decomposing images into editable layers through detection, segmentation and completion.
-While operationally sound, these multi-stage methods tend to accumulate errors due to their sequential processing pipelines, primarily manifested in two aspects: challenges in generating high-quality alpha channels for foreground layers and inconsistencies arising from amodal completion operations.
-This limitation catalyzed the emergence of direct RGBA synthesis paradigms.
-Text2Layer (Zhang et al., 2023b) marked a pivotal shift by jointly generating foreground/background layers through coordinated latent space optimization in Stable Diffusion.
-LayerDiffuse’s (Zhang and Agrawala, 2024) ”latent transparency” mechanism accelerated the field, which enabled pretrained LDMs to produce high-fidelity single transparent layers, and Alfie’s (Quattrini et al., [n.
-d.]) diffusion transformer architecture that dynamically derived alpha channels from attention maps.
+This limitation of pretrained generative models catalyzed the emergence of RGBA image layer synthesis.
+Text2Layer [Text2Layer] explored generating image layers through segmentation guidance.
+LayerDiffuse’s [zhang2024transparent] “latent transparency” mechanism accelerated the field, which enabled pretrained LDMs to produce high-fidelity single transparent layers.
 Building on transparent single-layer image generation, recent works have made strides toward multi-layer synthesis.
-LayerDiffuse (Zhang and Agrawala, 2024) itself extended its architecture for sequential background-to-foreground generation .
-Fontanella et al.
-(Fontanella et al., [n.
-d.]) explored independent multiple foreground layer generation and post-composition.
-LayerDiff (Huang et al., 2024) attempted to generate multi-layer images simultaneously, but it can only generate isolated, non-overlapping layers with binary masks.
+LayerDiffuse [zhang2024transparent] itself extended its architecture for sequential background-to-foreground generation .
+Fontanella et al[fontanellagenerating] explored separate multiple foreground layers generation and post-composition.
+LayerDiff [huang2024layerdiff] attempted to generate multi-layer images simultaneously, but it can only generate isolated, non-overlapping layers with binary masks.
 These paradigms mainly depend on straightforward layer stacking instead of considering multiple layers in a holistic manner.
@@
 Another critical challenge underlying this mult-layer image generation stems from the scarcity of high-quality multi-layer RGBA image datasets.
-While text-image pairs abound in datasets like LAION (Schuhmann et al., 2021), commercially restricted access to premium transparent assets severely limits usable multi-layer RGBA data.
+While text-image pairs abound in datasets like LAION [schuhmann2021laion400mopendatasetclipfiltered], commercially restricted access to premium transparent assets severely limits usable multi-layer RGBA data.
 Existing open-source multi-layer datasets exhibit complementary weaknesses.
-MAGICK (Burgert et al., 2024) and Multi-layer Dataset (Zhang and Agrawala, 2024) provide visually coherent layers but are constrained to 1-2 layers per composition, whereas MuLAn (Tudosiu et al., 2024) supports 2-6 layers per image but suffers from poor alpha-matte quality because it relies on automated post-processing pipelines.
+MAGICK [Burgert_2024_CVPR] and Multi-layer Dataset [zhang2024transparent] provide high quality layers but are constrained to 1-2 layers per composition, whereas MuLAn [tudosiu2024mulan] supports 2-6 layers per image but suffers from poor alpha-matte quality because it relies on automated post-processing pipelines.
 To overcome these dataset limitations, we propose the Inter-Layer Dataset, which consists of 30K meticulously curated multi-layer images, each containing 3-6 layers with artist-grade alpha mattes and rich layer interactions.
-Recognizing that automated synthesis and post-editing inherently fail to guarantee precise alpha channels and contextually coherent inpainting, we engaged professional designers to construct this dataset through a human-in-the-loop workflow.
-Designers curated layers from diverse sources including web repositories, BG20K (Li et al., 2022), MAGICK (Burgert et al., 2024), and LayerDiffuse-generated transparent assets (Zhang and Agrawala, 2024), and then applied meticulous edits to ensure natural layer interactions.
-Each composition integrates spatial layout optimization, physics-aware interaction and global stylistic harmony.
+Recognizing that automated synthesis and post-editing inherently fail to guarantee precise alpha channels and contextually coherent inpainting, we engaged about 35 professional designers to construct this dataset through a human-in-the-loop image composition workflow.
+Each composition operation involves spatial layout optimization, physics-aware layer interactions, and global stylistic consistency.
 Although existing pretrained text-to-image generation models can only generate unified single layer images, given a global text prompt describing the background, multiple foregrounds and their interactions, they did demonstrate prominent capability to spontaneously arrange the layout of multiple elements, ensure natural interactions between constituent subjects, and maintain global harmony and consistency in generated images.
-We believe that these pretrained text-to-image generation can provide interaction priors for multi-layer generation.
+We believe that these pretrained text-to-image generation models can provide interaction priors for multi-layer generation.
 Building on this fact, we propose PSDiffusion, a unified diffusion framework that utilizes the interaction priors from pretrained diffusion models, to simultaneously generate multi-layer images with coherent global structure and realistic layer interactions.
 We design an attention-based global-layer interactive mechanism to model the layer interactions paradigm.
-For one thing, to achieve a harmonious global layout for foreground layers, we develop a layer cross-attention reweighting module, where we extract the cross-attention map from global layer to guide the position of foreground layers.
+For one thing, to achieve a harmonious global layout for foreground layers, we develop a layer cross-attention reweighting module, where we extract the cross-attention map from the global layer to guide the position of foreground layers.
 For another, to facilitate content coherence and interaction among different layers, and avoid layer entanglement, we develop a partial joint self-attention module, where we encourage content sharing from global layer across the foreground and background layers.
@@
 Leveraging a global-layer interactive mechanism, our model generates layered images concurrently and collaboratively, ensuring not only high quality and completeness for each layer, but also spatial and visual interactions among layers for global coherence.
-2. Related Works
-2.1. Text-to-Image Generation
+Figure 2: Visual results of PSDiffusion.
+Through the design of Layer Cross-Attention Reweighting Module and Partial Joint Self-Attention Mechanism, PSDiffusion synthesize multi-layer transparent images with plausible layout arrangements and harmonious appearance.
+2 Related Work
+2.1 Text-to-Image Generation
 Diffusion models have made significant advancements in image generation, particularly in text-guided image synthesis, leading to remarkable improvements in both the quality of generated images and the ability to follow complex prompts.
-The development of large-scale diffusion models trained on extensive datasets of paired text and images has established new benchmarks in this domain, with notable examples including Stable Diffusion (Rombach et al., 2022), SDXL (Podell et al., 2024), Imagen (Saharia et al., 2022), and DALL-E 3 (Betker et al., 2023).
-Furthermore, recent advancements in network architecture, especially with Diffusion Transformer networks (DiT) like Stable Diffusion 3 (Esser et al., 2024) and Flux (Labs, 2024) , have enhanced image fidelity and diversity.
+The development of large-scale diffusion models trained on extensive datasets of paired text and images has established new benchmarks in this domain, with notable examples including Stable Diffusion [rombach2022high], SDXL [podell2023sdxl], Imagen [imagen], and DALL-E 3 [dalle3].
+Furthermore, recent advancements in network architecture, especially with Diffusion Transformer networks (DiT) like Stable Diffusion 3 [esser2024scaling] and Flux [flux2024] , have enhanced image fidelity and diversity.
 Moreover, other research efforts focus on integrating various control conditions into pre-trained image foundation models to enhance their controllability.
-For instance, Gligen (Li et al., 2023), ControlNet (Zhang et al., 2023a) and IP-Adapter (Ye et al., 2023) offer greater flexibility in controllable image generation, allowing users to precisely manipulate aspects such as layout, depth, human pose, and style.
+For instance, Gligen [li2023gligen], ControlNet [zhang2023control] and IP-Adapter [ye2023ip] offer greater flexibility in controllable image generation, allowing users to precisely manipulate aspects such as layout, depth, human pose, and style.
 However, these models can only generate a single RGB image during a single forward pass, rather than producing layered images, which are widely used formats that enable users to edit specific parts of an image without affecting other areas.
-2.2. Layered Image Generation
+2.2 Layered Image Generation
 Image Layer Extraction with Post-RGB Process.
 Multi-layer image generation can be achieved primarily through two different paths.
-The first path, termed Post-RGB Process, operates through a sequential pipeline: (1) initial RGB image synthesis via diffusion models (Rombach et al., 2022); (2) foreground extraction using open-vocabulary object detection and segmentation tools (e.g.
-GroundingDINO (Liu et al., 2024b), SAM (Kirillov et al., 2023), Matting-Anything (Li et al., 2024)); (3) background inpainting and foreground completion for occluded areas to derive multiple RGBA layers (Rombach et al., 2022; Podell et al., 2024; Yu et al., 2023; Ao et al., 2024).
-Along this way, Tudosiuet et al.
-(Tudosiu et al., 2024) developed a training-free pipeline, which decomposes a monocular RGB image into a stack of RGBA layers comprising background and isolated instances, and released the MuLAn dataset.
-LayeringDiff  (Kang et al., 2025) begins by generating using an off-the-shelf image generative model, followed by disassembling the image into its constituent foreground and background layers.
+The first path, termed Post-RGB Process, operates through a sequential pipeline: (1) initial RGB image synthesis via diffusion models [rombach2022high]; (2) foreground extraction using open-vocabulary object detection and segmentation tools (e.g.
+GroundingDINO [liu2024grounding], SAM [Kirillov2023Segment], Matting-Anything [li2023matting]); (3) background inpainting and foreground completion for occluded areas to derive multiple RGBA layers [rombach2022high, podell2023sdxl, yu2023inpaint, ao2024open].
+Along this way, Tudosiuet et al[tudosiu2024mulan] developed a training-free pipeline, which decomposes a monocular RGB image into a stack of RGBA layers comprising background and isolated instances, and released the MuLAn dataset.
+LayeringDiff [kang2025layeringdiff] begins by generating using an off-the-shelf image generative model, followed by disassembling the image into its constituent foreground and background layers.
 Yang et al.
-proposed LayerDecomp (Yang et al., 2024), a generative training framework that leverages both simulated and real-world data to enable robust layer-wise decomposition with accurate visual effect representation through a consistency loss, facilitating high-quality, training-free downstream editing.
-Generative Omnimatte (Lee et al., 2024) extends this idea to decompose video into layers and demonstrate high-quality decompositions and editing results.
-Despite the operational validity of these methods, the multi-stage post-processing approach imposes a nontrivial computational redundancy and risks compounding errors through iterative operations—a limitation that frequently manifests itself as incoherent visual semantics or fragmented stylistic attributes.
+proposed LayerDecomp [yang_Layer_Decomposition], a generative training framework that leverages both simulated and real-world data to enable layer-wise post-decomposition with visual effect representation.
+Despite the operational validity of these methods, the multi-stage post-processing approach imposes a nontrivial computational redundancy and risks compounding errors through iterative operations—a limitation that frequently manifests itself as incoherent visual semantics or poor alpha quality.
 Direct Transparent Image Layer Generation.
 RGBA direct generation focuses on synthesizing image layers directly, bypassing the traditional pipeline of generating a full image first and then extracting transparent layers.
-The earliest work, Text2Layer (Zhang et al., 2023b), extends the Stable Diffusion architecture into a dual-layer generation framework.
+The earliest work, Text2Layer [Text2Layer], extends the Stable Diffusion architecture into a dual-layer generation framework.
 It jointly synthesizes a background scene and semantically aligned foreground elements through coordinated latent space optimization.
-Subsequently, LayerDiffuse (Zhang and Agrawala, 2024) introduces ”latent transparency”, enabling large-scale pretrained latent diffusion models to generate either single transparent images.
+Subsequently, LayerDiffuse [zhang2024transparent] introduces ”latent transparency”, enabling large-scale pretrained latent diffusion models to generate either single transparent images.
 This approach has gained a reputation for its high-quality single-layer transparency generation.
-In parallel, Alfie (Quattrini et al., [n.
-d.]) and Diffumatting (Hu et al., 2024a) also develop different pipelines to generate single-layer RGBA images.
+In parallel, Alfie [quattrini2024alfie] and Diffumatting [hu2024diffumatting] also develop different pipelines to generate single-layer RGBA images.
 Building on single-layer generation capabilities, multi-layer synthesis is achieved through either sequential or separate generation paradigms.
-Specifically, LayerDiffuse (Zhang and Agrawala, 2024) adopts a sequential approach, proposing a background-conditioned transparent layer generation that iteratively synthesizes layers from background to foreground.
-Fontanella et al.
-(Fontanella et al., [n.
-d.]) and LayerFusion (Dalva et al., 2024) employ a separate paradigm, where individual foreground layers are generated independently and then composited.
-While methods of multi-layer generation based on single-layer generation have advanced RGBA synthesis, their extension suffers from insufficient layer interaction, leading to critical practical limitations.
-Common inter-layer contamination issues include rational global layout, physics-plausible contacts and visual effects like shadows and reflections.
-These problems are not easily resolvable through additional post-editing, particularly for the latter two issues.
-Recent unified text-guided multi-layer frameworks have emerged and shown promising results.
-Specifically, LayerDiff (Huang et al., 2024) proposes a simultaneous multi-layer synthesis framework using a unified diffusion model.
+Specifically, LayerDiffuse [zhang2024transparent] adopts a sequential approach, proposing a background-conditioned transparent layer generation that iteratively synthesizes layers from background to foreground.
+Fontanella et al[fontanellagenerating] and LayerFusion [dalva2024layerfusion] employ a separate paradigm, where individual foreground layers are generated independently and then composited.
+Recently unified text-guided multi-layer generation has attracted increasing attention.
+Specifically, LayerDiff [huang2024layerdiff] proposes a simultaneous multi-layer synthesis framework using a unified diffusion model with 3D convolutions.
 However, it produces low-quality outputs limited to binary masks instead of full alpha channels, failing to handle overlapping foregrounds and visual effects which are essential for layer editing.
-ART (Pu et al., 2025) introduces an Anonymous Region Layout Planner, aligning visual tokens with text tokens to directly synthesize spatially coherent multi-layer images.
-While innovative, it relies on precise bounding boxes provided by users or generated by LLMs.
-Additionally, its anonymous layout mechanism results in layer entanglement issue, where semantically distinct elements (e.g., foreground characters and background props) are incorrectly merged into one layer, adding unnecessary complexity.
-In contrast, our proposed PSDiffusion is a unified diffusion framework for simultaneous text to multi-layer image generation, ensuring not only high quality and completeness for each layer, but also spatial and semantic interaction among layers for global coherence.
-We adopt global prompt and layer-specific prompts as input to facilitate customized element separation and avoid layer entanglement.
-Figure 2.
-Architecture of PSDiffusion.
+ART [pu2025art] introduces an Anonymous Region Layout Planner, aligning visual tokens with text tokens to directly synthesize spatially coherent multi-layer images.
+However, it relies on precise bounding boxes provided by users or generated by LLMs and its anonymous layout mechanism results in layer entanglement issue, where semantically distinct elements (e.g., foreground characters and background props) are incorrectly merged into one layer, and cause issues on alpha quality.
+Figure 3: Architecture of PSDiffusion.
 Our framework processes multi-layer compositions through three key components: (1) A transparent VAE encoder preserving alpha channels; (2) Layer cross-attention reweighting for layout plausibility; (3) Partial joint self-attention for inter-layer context modeling.
-3. Method
-3.1. Problem Formulation
-A multi-layer composable image is defined as a structure containing a background layer $B=I_{c}^{B}\in\mathbb{R}^{H\times W\times 3}$ (RGB format) and $K$ foreground layers $\{F_{k}\}_{k=1}^{K}$ in RGBA format, where each $F_{k}=\left(I_{c}^{F_{k}},I_{\alpha}^{F_{k}}\right)$ consists of color channels $I_{c}^{F_{k}}\in\mathbb{R}^{H\times W\times 3}$ and an alpha channel $I_{\alpha}^{F_{k}}\in\mathbb{R}^{H\times W\times 1}$.
+3 Methodology
+3.1 Problem Formulation
+A multi-layer composable image is defined as a structure containing a background layer $B=I_{c}^{B}\in\mathbb{R}^{H\times W\times 3}$ in RGB format and $K$ foreground layers $\{F_{k}\}_{k=1}^{K}$ in RGBA format, where each $F_{k}=\left(I_{c}^{F_{k}},I_{\alpha}^{F_{k}}\right)$ consists of color channels $I_{c}^{F_{k}}\in\mathbb{R}^{H\times W\times 3}$ and an alpha channel $I_{\alpha}^{F_{k}}\in\mathbb{R}^{H\times W\times 1}$.
 Following the premultiplied alpha convention, each foreground is preprocessed as $F^{\prime}_{k}=I_{c}^{F_{k}}\ast I_{\alpha}^{F_{k}}$ through element-wise multiplication.
 The composition pipeline initializes with the background $\smash{I_{0}=I_{c}^{B}}$, then recursively blends layers via alpha operations. The $k$-th composite image is given by
-(1) $I_{k}=\left(1-I_{\alpha}^{F_{k}}\right)\cdot I_{k-1}+I_{\alpha}^{F_{k}}\ast I_{c}^{F_{k}},\ k\in[K]$
-where $F_{k}$ denotes the $k$-th foreground. After $K$ iterations, the final image $I_{K}$ composes into
-(2) $I_{K}=\underbrace{\left(\prod_{k=1}^{K}\left(1-I_{\alpha}^{F_{k}}\right)\right)\ast I_{c}^{B}}_{\text{\scriptsize Background visibility attenuated}}+\underbrace{\sum_{k=1}^{K}\left(\prod_{j=k+1}^{K}\left(1-I_{\alpha}^{F_{j}}\right)\ast I_{\alpha}^{F_{k}}\ast I_{c}^{F_{k}}\right)}_{\text{\scriptsize Foreground contributions with occlusion}}$
+$I_{k}=\left(1-I_{\alpha}^{F_{k}}\right)\cdot I_{k-1}+I_{\alpha}^{F_{k}}\ast I_{c}^{F_{k}},\ k\in[K]$ (1)
+where $F_{k}$ denotes the $k$-th foreground.
+The final composite image $I_{K}$ is gained after $K$ iterations.
 This mathematical formulation explicitly captures two fundamental mechanisms: (1) The multiplicative attenuation of background visibility through successive transparency operations; (2) The nested modulation of foreground layer contributions based on depth-ordered occlusions.
-3.2. PSDiffusion
-Existing multi-layer image generation models contend with inherent challenges in synthesizing coherent compositions and fostering harmonious interactions between distinct high-quality transparent foreground layers and the background image (Tudosiu et al., 2024; Zhang and Agrawala, 2024).
+3.2 PSDiffusion
+Existing multi-layer image generation models contend with inherent challenges in synthesizing coherent compositions and fostering harmonious interactions between distinct high-quality transparent foreground layers and the background image [tudosiu2024mulan, zhang2024transparent].
 However, it has been observed that pretrained image foundation models demonstrate proficiency in generating intricate and visually coherent compositions, effectively integrating various objects and backgrounds within a single RGB image, owing to their training on extensive datasets derived from real-world imagery.
-Our primary insight is to exploit the layout and interaction priors from the denoising process of the global RGB image, to inform the generation of transparent foreground layers and background layer.
-To fully take advantage of the layout arrangement and appearance harmonization of the global RGB images, we propose a novel framework, denoted as PSDiffusion, which concurrently generates multiple layers in a single feedforward pass, ensuring both reasonable layouts and harmonized interactions among the layers.
+Our primary insight is to exploit the layout and interaction priors from the denoising process of the global RGB image, to guide the generation of foreground and background layers.
+To fully take advantage of the layout arrangement and appearance harmonization of the global RGB models, we propose a novel framework, denoted as PSDiffusion, which concurrently generates multiple layers in a single feedforward pass, ensuring both reasonable layouts and harmonized interactions among the layers.
 As delineated in Figure 3, PSDiffusion comprises three principal components.
-First, PSDiffusion finetunes a pretrained RGB image Variational Autoencoder (VAE) into an RGBA VAE, thereby facilitating the generation of the alpha channel requisite for RGBA images, as articulated by LayerDiffuse (Zhang and Agrawala, 2024).
+First, PSDiffusion finetunes a pretrained RGB image Variational Autoencoder (VAE) into an RGBA VAE, thereby facilitating the generation of the alpha channel requisite for RGBA images, as articulated by LayerDiffuse [zhang2024transparent].
 Second, we introduce a layer cross-attention reweighting mechanism that extracts spatial layout information from the global image containing comprehensive content, generated via a global prompting strategy.
@@
 Finally, to promote harmonious interactions among the various foreground layers, we implement a partial joint self-attention module for inter-layer context modeling.
-Detailed descriptions of these two modules are provided as follows.
-Figure 3. The overview of layer cross-attention reweighting module and partial joint self-attention mechanism.
+Figure 4: Overview of our global-layer interactive mechanism, composed of layer cross-attention reweighting module and partial joint self-attention module.
+Layer cross-attention reweighting module extracts the cross-attention map from the global branch to reweight the cross-attention map of the layer branch, guiding the position of foreground layers.
+Partial joint self-attention module implements a shared attention across global branch and layer branch to facilitate context-aware feature modeling.
 Layer Cross-Attention Reweighting.
@@
 As illustrated in Figure 4, we utilize the cross-attention block of pre-trained image foundation model to extract the layout priors.
-Leveraging the emergent correspondence between prompts and features—a phenomenon widely demonstrated and utilized in existing literature (Hertz et al., [n.
-d.]; Tang et al., 2023; Cao et al., 2023; Epstein et al., 2023; Hu et al., 2024b)—we can precisely derive and control the location of foreground instance by calculating and manipulating the cross-attention map associated with the foreground instance token.
-As is highlighted in the prompts in Figure 3, we select the token of the noun representing the foreground subject because its attention map most prominently displays the location of the foreground subject.
-Specifically, at timestep $t$ in the $n$-th cross-attention block, the global text embedding $\psi(p)$ is projected to a key matrix $\mathbf{K}^{n}=\ell_{\mathbf{K}}^{n}(\psi(p))$ and a value matrix $\mathbf{V}^{n}=\ell_{\mathbf{V}}^{n}(\psi(p))$ , and the deep spatial features of the noisy image $\phi(z_{t})$ is projected to a query matrix $\mathbf{Q}^{n}=\ell_{\mathbf{Q}}^{n}(\phi(z^{t}))$, via learned linear projections $\ell_{\mathbf{K}}^{n}$, $\ell_{\mathbf{V}}^{n}$, and $\ell_{\mathbf{Q}}^{n}$, respectively.
+Leveraging the emergent correspondence between prompts and features—a phenomenon widely demonstrated and utilized in existing literature [hertzprompt, tang2023emergent, cao_2023_masactrl, epstein2023diffusion, hu2023anomalydiffusion]—we can precisely derive and control the location of foreground instance by calculating and manipulating the cross-attention map associated with the foreground instance token.
+As shown in Figure 3, we select the token of the colored noun representing the foreground subject because its attention map most prominently displays the location of the foreground subject.
+At timestep $t$ in the $n$-th cross-attention block, the global text embedding $\psi(p)$ is projected to a key matrix $\mathbf{K}^{n}=\ell_{\mathbf{K}}^{n}(\psi(p))$ and a value matrix $\mathbf{V}^{n}=\ell_{\mathbf{V}}^{n}(\psi(p))$ , and the deep spatial features of the noisy image $\phi(z_{t})$ is projected to a query matrix $\mathbf{Q}^{n}=\ell_{\mathbf{Q}}^{n}(\phi(z^{t}))$, via learned linear projections $\ell_{\mathbf{K}}^{n}$, $\ell_{\mathbf{V}}^{n}$, and $\ell_{\mathbf{Q}}^{n}$, respectively.
 The global attention map $\mathbf{M}^{\mathrm{g}}\in\mathbb{R}^{hw\times s}$can be calculated by
-(3) $\mathbf{M}^{\mathrm{g}}=\mathrm{Softmax}\left(\dfrac{\mathbf{Q}^{n}\left(\mathbf{K}^{n}\right)^{\mathrm{T}}}{\sqrt{d}}\right).$
+$\mathbf{M}^{\mathrm{g}}=\mathrm{Softmax}\left(\dfrac{\mathbf{Q}^{n}\left(\mathbf{K}^{n}\right)^{\mathrm{T}}}{\sqrt{d}}\right).$ (2)
 We extend our method to obtain the layer attention map $M^{l}\in\mathbb{R}^{hw\times s}$ in the latent feature of the transparent image layer using an analogous calculation approach, where $hw$ denotes the size of the latent feature and $s$ denotes the length of the tokens.
 We extract the specific attention-map of the token representing the foreground subject from the global map and layer map, denoted as $\mathbf{M}_{i}^{\mathrm{g}}\in\mathbb{R}^{hw}$ and $\mathbf{M}_{i}^{\mathrm{l}}\in\mathbb{R}^{hw}$.
-Then implement the layer cross attention map reweighting by
-(4) ${\tilde{\mathbf{M}}_{i}^{l}}$ $=\mathrm{Norm}\left(\mathbf{M}_{i}^{\mathrm{g}}\right)*\mathbf{M}_{i}^{l},$
-(5) ${\hat{\mathbf{M}}_{i}^{l}}$ $={\tilde{\mathbf{M}}_{i}^{l}}/\mathrm{max}\left(\tilde{\mathbf{M}}_{i}^{l}\right)*\mathrm{max}\left(\mathbf{M}_{i}^{l}\right),$
-where $\mathrm{Norm}(\cdot)$ denotes Min-Max normalization on the whole feature dimension and $\max(\cdot)$ is applied on the feature dimension as well.
+Since there are differences between the global map $\mathbf{M}_{i}^{\mathrm{g}}$ and layer map $\mathbf{M}_{i}^{\mathrm{l}}$ during the original denoising process, we utilize the global map as a guide to reweight the layer map.
+We implement this layer cross attention map reweighting by
+$\tilde{\mathbf{M}}_{i}^{\mathrm{l}}=\mathrm{Norm}\left(\mathbf{M}_{i}^{\mathrm{g}}\right)\odot\mathbf{M}_{i}^{\mathrm{l}},$ (3)
+$\hat{\mathbf{M}}_{i}^{\mathrm{l}}=\tilde{\mathbf{M}}_{i}^{\mathrm{l}}\cdot\dfrac{\max\left(\mathbf{M}_{i}^{\mathrm{l}}\right)}{\max\left(\tilde{\mathbf{M}}_{i}^{\mathrm{l}}\right)}.$ (4)
+where $\mathrm{Norm}(\cdot)$ denotes Min-Max normalization on the whole feature dimension, $\odot$ denotes Hadamard Product, $\max(\cdot)$ is applied on the feature dimension as well for stability,.
 This process provides a more precise attention map that clearly delineates the position and shape of areas associated with the foreground layer prompt.
-The refined attention mask is then strategically integrated into the self-attention blocks of the latent diffusion model.
+The global attention map is then strategically processed and integrated into the self-attention blocks of the latent diffusion model.
 Through this approach, we achieve a more nuanced and controlled layout arrangement for multi-layer image generation.
 Partial Joint Self-attention.
-While our proposed layer cross-attention reweighting mechanism ensures a reasonable layout arrangement for each layer, challenges remain in harmonizing the color, style and interactions across different layers.
+While our proposed layer cross-attention reweighting mechanism ensures a reasonable layout arrangement for each layer, challenges remain in harmonizing the color, style and physics interactions across different layers.
 To address the appearance inconsistency among layers, we introduce a novel partial joint self-attention mechanism for cross-layer feature context modeling.
-This approach facilitates more nuanced interactions among layers, ultimately achieving consistent color and lighting conditions within the composite image.
-As illustrated in Figure 3, we concatenate the features from global branch into the $n$-th self-attention block of the layer branch,
-(6) $\mathbf{Q}=\mathbf{Q}_{L}=\mathbf{W}^{n}_{\mathbf{Q}}z_{i},\ \mathbf{K}=\mathbf{W}^{n}_{\mathbf{K}}\left[\mathbf{K}_{GL},\mathbf{K}_{L}\right],\ \mathbf{V}=\mathbf{W}^{n}_{\mathbf{V}}\left[\mathbf{V}_{GL},\mathbf{V}_{L}\right],$
-where $z_{i}$ denotes the latent feature for the $i$-th transparent layer, [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes the concatenation operation. $\mathbf{K}_{GL}$ and $\mathbf{V}_{GL}$ is calculated by
-(7) $\mathbf{K}_{GL}=\{\mathbf{K}_{G}\cdot\mathrm{Mask}_{i}^{\mathrm{g}}\},\ \mathbf{V}_{GL}=\{\mathbf{V}_{G}\cdot\mathrm{Mask}_{i}^{\mathrm{g}}\}.$
+This approach facilitates more nuanced interactions among layers, ultimately achieving consistent color, geometry and lighting conditions within the composite image.
+As illustrated in Figure 4, we concatenate the features of the global branch into the $n$-th self-attention block of the layer branch,
+$\mathbf{Q}=\mathbf{Q}_{L}=\mathbf{W}^{n}_{\mathbf{Q}}z_{i}$ (5)
+$\mathbf{K}=\mathbf{W}^{n}_{\mathbf{K}}\left[\mathbf{K}_{GL},\mathbf{K}_{L}\right],\ \mathbf{V}=\mathbf{W}^{n}_{\mathbf{V}}\left[\mathbf{V}_{GL},\mathbf{V}_{L}\right]$ (6)
+where $z_{i}$ denotes the latent feature for the $i$-th transparent layer, $\cdot$$\cdot$] denotes the concatenation operation. $\mathbf{K}_{GL}$ and $\mathbf{V}_{GL}$ are calculated by
+$\mathbf{K}_{GL}=\{\mathbf{K}_{G}\cdot\mathrm{Mask}_{i}^{\mathrm{g}}\},\ \mathbf{V}_{GL}=\{\mathbf{V}_{G}\cdot\mathrm{Mask}_{i}^{\mathrm{g}}\}.$ (7)
+where $\mathrm{Mask}_{i}^{\mathrm{g}}$ is derived from the global attention map $\mathbf{M}_{i}^{\mathrm{g}}$ in the Layer Cross-Attention Reweighting block and applied with filtering, blurring, and binarization, $\{\cdot\}$ denotes that we apply the attention mask in the multi-head attention.
+With these masks, we can restrict the $i$-th layer to query context information only from the corresponding and nearby regions in the global image, thus avoiding layer entanglement issues.
+Meanwhile, due to the inherent appearance harmony and element interactions of the global image, the content sharing from the associate regions of global branch to layer branch is enough for the composite harmonization.
+Overall, such a design enables the model to refer to the appearance features of the global layer, promoting a more cohesive and harmonized image generation across different foreground transparent layers and background layer.
+Figure 5: Examples of our proposed Inter-Layer dataset.
+3.3 Inter-Layer Dataset
+To address the limitations of existing multi-layer datasets, we developed the dataset Inter-Layer through a human-centric curation pipeline.
+This dataset comprises 30K high-fidelity multi-layer compositions, each containing 3 - 6 layers with professional-grade alpha mattes and sophisticated inter-layer relationships.
+Recognizing the inadequacy of automated synthesis for achieving precise transparency channels and contextual coherence, we engaged professional designers in a multi-stage workflow.
+Curators systematically curated assets from diverse sources including web repositories, BG20K [li2022bridging], MAGICK [Burgert_2024_CVPR], LayerDiffuse-generated transparent assets [zhang2024transparent] and Flux-generated assets [flux2024], followed by rigorous manual refinement.
+Designers optimized each composition through: (1) Spatial layout organization with an aesthetic consideration; (2) Physics-plausible interactions among all layers, especially geometry and orientation consistency.
+Designers strategically select layer combinations from our extensive RGBA asset library whose spatial and geometry relationships are pre-aligned.
+Minor 2D rotational and shape adjustments are then applied to optimize spatial coherence and guarantee physics-plausible contacts.
+(3) Global visual harmonization of all elements.
+Designers perform comprehensive visual adjustments to maintain consistency between foreground objects and their contextual environment.
+This includes calibrating saturation, hue, contrast, resolution, sharpness and brightness to match surrounding conditions.
+Additionally, shadows and reflections are manually crafted according to environmental factors (e.g., water surfaces, lighting direction) to achieve photorealistic integration.
+This human-in-the-loop approach ensures both technical precision in alpha channel quality and visual consistency in multi-layer storytelling, significantly advancing beyond existing datasets’ capabilities in layer complexity and visual authenticity.
+Figure 5 presents examples from our proposed dataset, highlighting its high quality, diverse multi-foreground layers, harmonized compositions, and consistent appearance.
+Figure 6: Qualitative comparison with state-of-the-art methods.
+Compared to existing methods, our model synthesizes multi-layer images with reasonable layout arrangement, harmonious appearance with visual effects and high alpha quality.
+Zoom in to check the details of foreground and composite images like shadows, reflections and foreground edges.
+As exemplified by the multi-layer ensemble in the upper-right quadrant, the shadows of the elephant, man and woman are consistent with the background in their respective positions.
+4 Experiments
+4.1 Implementation Details
+To better preserve the interaction priors of existing models, PSDiffusion is fine-tuned from Stable Diffusion XL [podell2023sdxl] using LoRA [hu2022lora].
+For global layer, we trained the additional $\mathbf{K}_{GL}$$\mathbf{V}_{GL}$ linear layers in all self-attention blocks for joint attention.
+For foreground layers and background layers, we trained $\mathbf{Q}$$\mathbf{K}$$\mathbf{V}$$\mathbf{Out}$ linear layers in all attention blocks.
+Layer Cross-Attention Reweighting is implemented in layers at the resolution of 16, and the Partial Joint Self-attention is applied to all self-attention layers.
+LoRA rank is set to 256.
+We use the pre-trained VAE from Stable Diffusion XL for the global and background branch, and fine-tune it into an RGBA VAE to support the alpha channel of RGBA foreground image following LayerDiffuse [zhang2024transparent].
+During training, all layers of a multi-layer sample are encoded into latent space and added with the same noise at the same timestep.
+We use the AdamW optimizer and set the learning rate at $1e\text{-}4$.
+Training is carried out over our Inter-Layer dataset with a total batch size of 8 on 60 NVIDIA A800 GPU hours for 10,000 iterations.
+4.2 Compared with State-of-the-Art
+Method Composite Image Layer Image
 […]
-Overall, such a design enables the model to refer the appearance of the global layer, promoting a more cohesive and harmonized image generation across different foreground transparent layers and background layer.
-Figure 4. Examples of our proposed inter-layer dataset.
-3.3. Inter-Layer Dataset
-To address the limitations of existing multi-layer datasets, we developed the dataset Inter-Layer through a human-centric curation pipeline.
-This dataset comprises 30,000 high-fidelity multi-layer compositions, each containing 3 - 6 layers with professional-grade alpha mattes and sophisticated inter-layer relationships.
-Recognizing the inadequacy of automated synthesis for achieving precise transparency channels and contextual coherence, we engaged professional designers in a multi-stage workflow.
-Curators systematically aggregated assets from diverse sources including web repositories, BG20K (Li et al., 2022), MAGICK (Burgert et al., 2024), and LayerDiffuse-generated transparent assets (Zhang and Agrawala, 2024), followed by rigorous manual refinement.
-Designers optimized each composition through: (1) Spatial layout organization with an aesthetic consideration; (2) Physics-plausible interactions among all layers; (3) Global visual harmonization across all elements.
-This human-in-the-loop approach ensures both technical precision in alpha channel quality and visual consistency in multi-layer storytelling, significantly advancing beyond existing datasets’ capabilities in layer complexity and visual authenticity.
-Some figures in our dataset are shown in Figure 4.
-Figure 5.
-Qualitative comparison with state-of-the-art methods.
-Compared to existing methods, our model synthesizes multi-layer images with reasonable layout arrangement, harmonious appearance with visual effects and high alpha quality.
-Zoom in to check the details of foreground and composite images like shadows and edges.
-As exemplified by the multi-layer ensemble in the upper-right quadrant, the shadows of the elephant, man and woman are consistent with the background in their respective positions.
-4. Experiments
-4.1. Implementation Details
-To better preserve the interaction priors of existing models, PSDiffusion is fine-tuned from Stable Diffusion XL (Podell et al., 2024) using LoRA (Hu et al., 2022).
-For global layer, we trained the additional $\mathbf{K}_{GL}$,$\mathbf{V}_{GL}$ linear layers in all self-attention blocks.
-For foreground layers and background layers, we trained $\mathbf{Q}$,$\mathbf{K}$,$\mathbf{V}$,$\mathbf{Out}$ linear layers in all attention blocks.
-Layer Cross-Attention Reweighting is implemented in layers at the resolution of 16, and the Partial Joint Self-attention is applied to all self-attention layers.
-We use the pre-trained VAE from Stable Diffusion XL for the global and background branch, and fine-tune it into an RGBA VAE to support the alpha channel of RGBA foreground image following LayerDiffuse (Zhang and Agrawala, 2024).
-During training, all layers of a multi-layer sample are encoded into latent space and added with the same noise at the same timestep.
-We use the AdamW optimizer and set the learning rate at $-41e\text{-}$ 4.
-Training is conducted with a total batch size of 8 on 20 NVIDIA A800 GPU hours for 10,000 iterations.
-4.2. Compared with State-of-the-Art
-Composite Image Layer Image
-[…]
-LayerDiffuse  (Zhang and Agrawala, 2024) 29.77 89.12 0.265 0.214 31.25 85.85
-ART  (Pu et al., 2025) 29.93 96.54 0.743 0.702 128.1
+LayerDiffuse [zhang2024transparent] 29.77 89.12 0.265 0.214 31.25 85.85
+ART [pu2025art] 29.93 96.54 0.743 0.702 128.1
 PSDiffusion(ours) 31.89 87.32 0.766 0.751 31.76 83.71
-Table 1.
-Quantitative comparison with state-of-the-art methods.
+Table 1: Quantitative comparison with state-of-the-art methods.
 We report performance on both composite images and layer images in terms of visual quality, text alignment and composite layout harmony and interaction plausibility.
-Figure 6.
-More visual results of PSDiffusion.
-Through the design of Layer Cross-Attention Reweighting Module and Partial Joint Self-Attention Mechanism, PSDiffusion synthesize multi-layer transparent images with plausible layout arrangements and harmonious appearance.
-We compare the quality of multi-layer images generated by our model with LayerDiffuse (Zhang and Agrawala, 2024) and the latest work ART (Pu et al., 2025), as they are the state-of-the-art layered image synthesis methods, and the only methods that release the source code.
+We compare our model with LayerDiffuse [zhang2024transparent] and the latest work ART [pu2025art], as they are the state-of-the-art layered image synthesis methods, and the only methods that release the source code.
 For evaluation, we construct 5,000 sets of multi-layer prompts, each composed of three foreground prompts, one background prompt and a global prompt describing their spatial and semantic interactions.
-We use vision-language model (Bai et al., 2025) to generate the prompts to establish diverse evaluation scenarios.
+We use vision-language model [Qwen2.5-VL] to generate the prompts to establish diverse evaluation scenarios.
 We conduct quantitative evaluation using multiple metrics.
-For image quality assessment, we measure feature distribution discrepancies between synthesized layers and reference datasets via FID (Heusel et al., 2017) (Fréchet Inception Distance).
-Specifically, the RGBA layer is evaluated against the MAGICK dataset (Burgert et al., 2024), while the composite images are compared with the COCO dataset (Lin et al., 2014).
-To assess text-image alignment for each layer, we employ the CLIP Score metric (Hessel et al., 2021), which quantifies the cosine similarity between image features and text prompts.
-To evaluate the harmony of the global layout and the plausibility of the layer interactions, we utilize vision-language model (Bai et al., 2025) to conduct scoring.
+For image quality assessment, we measure feature distribution discrepancies between synthesized layers and reference datasets via FID [heusel2017gans] (Fréchet Inception Distance).
+Specifically, the RGBA layer is evaluated against the MAGICK dataset [Burgert_2024_CVPR], while the composite images are compared with the COCO dataset [lin2014microsoft].
+To assess text-image alignment for each layer, we employ the CLIP Score metric [hessel2021clipscore], which quantifies the similarity between image features and text prompts.
+To evaluate the harmony of the global layout and the plausibility of the layer interactions, we utilize vision-language model [Qwen2.5-VL] to conduct scoring comprehensively.
+The definition of layout harmony includes the rationality of the global layout and the scale of each element.
+The definition of interaction plausibility includes physical contacts and geometry consistency among entities, coordination of color and style, and consistency of light and shadows.
 Qualitative Results.
-In Figure  5, we present the qualitative results on text prompt multi-layer image generation.
-LayerDiffuse (Zhang and Agrawala, 2024) tends to stacks all layers at the center of the image, neglecting the spatial interactions among all layers.
+In Figure 6, we present the qualitative results on text prompt multi-layer image generation.
+LayerDiffuse [zhang2024transparent] tends to stack all layers at the center of the image, neglecting the spatial interactions among all layers.
 The latest work ART showcases reasonable layout but relies on the precise bounding boxes given by users.
-It also exihibits the problem of imperfect alpha quality and layer entanglement.
-In contrast, our model displaces ont only high quality RGBA layers, but also the capability to arrange the layers to reasonable scale and postion, and mainting the visual interactions among the layers like the semi-transparent shadows.
-Zoom in to see the details of foreground and composite images like edges and shadows.
+It also exhibits the problem of imperfect alpha quality and layer entanglement.
+In contrast, our model displays not only high quality of RGBA layers, but also the capability to arrange the layers to reasonable scale and position, and maintaining the visual interactions among the layers.
+Zoom in to see the details of foreground and composite images like edges, shadows and reflections.
+More qualitative results of our model are shown in Figure 2.
 Quantitative Results.
-We report quantitative results in Table  1.
-We compare our PSDiffusionwith the state-of-the-art multi-layer image generation models LayerDiffuse (Zhang and Agrawala, 2024) and ART (Pu et al., 2025) on composite images.
-For LayerDiffuse (Zhang and Agrawala, 2024), we implement the proposed sequential background-to-foreground geneartion to gain the multi-layer images.
-For ART (Pu et al., 2025), since it relys on the bounding-boxes as input, we first apply the LLM layout generation proposed by the authors given our global prompt, and then use the generated bounding boxes and global prompts to generate the multi-layer images.
+We report quantitative results in Table 1.
+We compare our PSDiffusionwith the state-of-the-art multi-layer image generation models LayerDiffuse [zhang2024transparent] and ART [pu2025art] on composite images.
+For LayerDiffuse, we implement the proposed sequential background-to-foreground generation to gain the multi-layer images.
+For ART, since it relies on the bounding-boxes as input, we first apply the LLM layout generation proposed by the authors given our global prompt, and then use the generated bounding boxes and global prompts to generate the multi-layer images.
 As shown in the table, our PSDiffusion outperforms LayerDiffuse and ART across four metrics.
 The results on layout harmony and interaction plausibility directly showcase our strengths on generating multi-layer image with enhanced layer interaction and global rationality.
-Furthermore, for the alpha image layer, the results in Table  1 shows that our PSDiffusion also achieves excellent alpha image quality and better text alignment.
+Furthermore, for the alpha image layer, the results in Table 1 shows that our PSDiffusion also achieves excellent alpha image quality and better text alignment.
 User Study.
@@
 Each sample underwent multi-criteria evaluation using 5-point Likert scales (1: deficient, 5: exceptional) focusing on three critical dimensions: (1) semantic-textual congruence measuring prompt relevance and contextual appropriateness, (2) visual-perceptual fidelity evaluating aesthetic coherence and artifact suppression, and (3) structural clarity assessing image definition and detail preservation.
-The comparative evaluation results are summarized in Table  2, demonstrating that our method consistently achieves higher user ratings across than other methods.
-Complete experimental details regarding the user study design and implementation are available in the supplementary materials.
+The comparative evaluation results are summarized in Table 2, demonstrating that our method consistently achieves higher user ratings across than other methods.
+4.3 Ablation Studies
+Layer Cross-Attention Reweighting.
+In this module, we utilize the cross-attention map of the global layer, and implement the reweighting operation on the cross attention map of the foreground layers to control the global layout.
+As shown in Figure 7, without the layer attention-map reweighting, the foreground object tends to occupy the most of the image like LayerDiffuse and even shows poor quality on the alpha channel, especially for the foreground that should have been small in the composite image.
+Figure 7: Visual results of ablation study for layer cross-attention reweighting. Full PSDiffusion exhibits more plausible spatial layout including postion and scale.
+Partial Joint Self-Attention.
+In our partial joint self-attention blocks, we implement the content sharing from global layer across foreground and background layers to encourage layer interaction plausibility like the shadows.
+Figure 8 displays that removing this block leads to visual effects like rigid stacks of separated foreground layers hence global disharmony.
+As shown in Table 3, removing either partial joint self-attention block or layer cross-attention reweighting block results in significant decline in both layout harmony and layer interactions.
+This showcases their entangled importance to the coherence and interaction plausibility of the composite image.
+Figure 8: Visual results of ablation study for partial joint self-attention. Full PSDiffusion demonstrates more coherent appearances and visual effects (e.g., shadows).
 […]
-LayerDiffuse (Zhang and Agrawala, 2024) 1.64 3.63 1.53
-ART (Pu et al., 2025) 1.30 2.53 3.14
+LayerDiffuse [zhang2024transparent] 1.64 3.63 1.53
+ART [pu2025art] 1.30 2.53 3.14
 PSDiffusion(ours) 2.06 3.83 3.35
-Table 2. User study on multi-layer image generation models. Results generated by our PSDiffusionare the most popular with users.
-4.3. Ablation studies
+Table 2: User study on multi-layer image generation models. Our PSDiffusion are the most popular with users.
 […]
 w/o Reweighting 31.64 89.18 0.565 0.418
-w/o Joint Attention 30.71 89.54 0.488 0.350
+w/o Joint Atten. 30.71 89.54 0.488 0.350
 Full model 31.89 87.32 0.766 0.751
-Table 3. Ablation study on Layer Cross-Attention Reweighting and Partial Joint Self-Attention on composite images
-Layer Cross-Attention Reweighting.
-In our layer cross-attention reweighting, we utilize the cross-attention map of the global layer, and implement the reweighting operation on the cross attention map of the foreground layers to control the global layout.
-As shown in Figure 7, without the layer attention-map reweighting, the foreground object tends to occupy the most of the image like LayerDiffuse and even shows poor quality on the alpha channel, especially for the foreground that should have been small in the composite image.
-Figure 7. Visual results of ablation study for layer cross-attention reweighting. Full PSDiffusion exihibits more plausible spatial layout including postion and scale.
-Partial Joint Self-Attention.
-In our partial joint self-attention blocks, we implement the content sharing from global layer across foreground and background layers to encourage layer interaction plausibility like the shadows.
-Figure  8 displays that removing this block leads to visual effects like rigid stacks of separated foreground layers hence global disharmony.
-As shown in Table  3, removing either Partial joint self-attention block or Layer cross-attention reweighting block results in significant decline in both global harmony and layer interactions.
-This showcases their entangled importance to the coherence and interaction plausibility of the composite image.
-Figure 8. Visual results of ablation study for partial joint self-attention. Full PSDiffusion demonstrates more coherent appearances and visual effects (e.g., shadows).
-5. Conclusion
+Table 3: Ablation study on Layer Cross-Attention Reweighting and Partial Joint Self-Attention on composite images
+5 Conclusion
 In this paper, we propose the Inter-Layer Dataset, a high-quality multi-layer RGBA image dataset with artist-grade alpha mattes and rich layer interactions.
@@
 Our model generates layered images concurrently and collaboratively, ensuring not only high quality and completeness for each layer, but also spatial and visual interactions among layers for global coherence.
-For one thing, to achieve a harmonious global layout for foreground layers, we develop a layer cross-attention reweighting module, which extracts the cross-attention maps from the global branch to guide the positioning of foreground layers.
-For another, to enable content coherence and interaction among different layers, and avoid layer entanglement, we develop a partial joint self-attention module, where we facilitate content sharing from global layer across the foreground and background layers.
-With these layout and appearance alignment, we empower diffusion models to generate multi-layer images with enhanced layer interactions and composite harmony.
-Supplementary Materials
-Appendix A Limitations
-A.1. Planar composition constraints in volumetric interaction modeling
-While our model achieves sophisticated occlusion handling through discrete 2D layer composition, its planar representation inherently limits the modeling of volumetric spatial interactions.
-The strict layer-wise depth prioritization (e.g., globally enforcing Layer A to occlude Layer B) cannot capture localized mutual occlusion in 3D space, such as intertwined limbs during human embraces.
-Static layer ordering further fails to adapt to perspective-dependent depth transitions under viewpoint changes, while non-planar intersections with depth-dependent partial visibility remain unresolved due to reliance on discrete alpha compositing.
-These constraints stem from our use of discretized depth ordering rather than continuous 3D occupancy field learning.
-A.2. Uniform sharpness artifacts from 2D-flattened rendering
-Despite achieving exceptional per-layer clarity and lighting consistency, the uniform maximum sharpness across all layers prevents emulation of natural depth-of-field effects (e.g., foreground-background focus gradients).
-This limitation arises from the absence of depth-aware dynamic blurring, compromising visual realism in scenes requiring strong 3D spatial perception.
-The persistent all-in-focus appearance introduces planar artifacts, particularly noticeable in multi-layer compositions with complex depth relationships, as the discrete compositing mechanism cannot simulate depth-dependent optical blur.
-Appendix B Dataset Visualization
-Benefiting from our human-centric curation pipeline and professional-grade multi-stage refinement, our dataset achieves three key superior qualities: high-fidelity alpha mattes, physics-aware inter-layer interactions, and globally harmonized visual aesthetics.
-Fig.
-910 are visualizations of our dataset compositions.
-Figure 9. Visualization of our multi-layer dataset
-Figure 10. More visulization of our multi-layer dataset
-Appendix C User study
-The following Table 4 presents the questionnaire form we made for the investigation in the userstudy part.
-Table 4. User Study of Multi-Layer Generation Models
-Current AI image generators create flat images, unlike professional design workflows using layered files (e.g., Photoshop .psd). Our new system automatically generates:
-multi-layer images (1 background + multiple RGBA layers) Context-aware layer coordination Production-ready layered files
-Evaluation Criteria
-1. Semantic-textual congruence
-Relevance to prompts & contextual logic Visual-perceptual fidelity 2. Aesthetic unity & artifact-free rendering
-3. Structural clarity
-Detail preservation & layer definition
-Your Task
-For each text-generated image:
-Assign a 1-5 overall score (1=worst, 5=best) Base your rating on the three criteria above
-Example Case
-”whole text”: A white cat sitting and looking at the camera, positioned in front of a wooden floor Foreground:A white cat [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Score Background:wooden floor [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Score Merged image [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Score
+Future Work.
+Given that our method leverages the intrinsic interaction priors of pre-trained generative models and achieves excellent multi-layer generation results through an attention-based global-layer interactive module, it holds the potential to be migrated to other attention-based pre-trained generative models.
+In the future, we plan to explore extending our method to a broader range of generative models, including DiT-based image generation [esser2024scaling, flux2024], video generation [wan2025, blattmann2023stable], and 3D generation models [lin2023magic3d].